# Two axes in non-associative algebras with a Frobenius form

### v5

@@ 5. The proof that $b_{\frac{1}{2}}(x_{0}y_{0})=(b_{\frac{1}{2}}x_{0})y_{0}+(b_{\frac{1}{2}}y_{0})x_{0}$
$-\frac{1}{2}(b,y_{0})(a,b)b_{\frac{1}{2}}x_{0}-(b,y_{0})(b_{\frac{1}{2}}x_{0})b_{0}-(b,b_{\frac{1}{2}}x_{0})b_{\frac{1}{2}}y_{0}-(b,(b_{\frac{1}{2}}x_{0})y_{0})b_{\frac{1}{2}}+((b_{\frac{1}{2}}x_{0})y_{0})b_{0}$
@@
$-\frac{1}{4}(b,x_{0})b_{\frac{1}{2}}y_{0}-\frac{1}{2}(b,y_{0})(a,b)b_{\frac{1}{2}}x_{0}-\frac{1}{4}(b,x_{0})(b,y_{0})b_{\frac{1}{2}}-(a,b)(b,x_{0})b_{\frac{1}{2}}y_{0}-(b,(b_{\frac{1}{2}}x_{0})y_{0})b_{\frac{1}{2}}$
@@
$(2(a,b)-1)(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})+(b,x_{0})\Big(\frac{1}{2}(1-(a,b))b_{\frac{1}{2}}y_{0}+\frac{1}{4}(b,y_{0})b_{\frac{1}{2}}\Big)+2(a,b)(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})$
$-\frac{1}{4}(b,x_{0})b_{\frac{1}{2}}y_{0}-\frac{1}{2}(b,y_{0})(a,b)b_{\frac{1}{2}}x_{0}-\frac{1}{4}(b,x_{0})(b,y_{0})b_{\frac{1}{2}}-(b,(b_{\frac{1}{2}}x_{0})y_{0})b_{\frac{1}{2}}$
@@
$2(a,b)(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})+2(a,b)(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})-\frac{1}{2}(a,b)(b,y_{0})b_{\frac{1}{2}}x_{0}-\frac{1}{2}(a,b)(b,x_{0})b_{\frac{1}{2}}y_{0}$
$-\frac{1}{2}(a,b)(x_{0}b_{0},y_{0})b_{\frac{1}{2}}$
$+\frac{1}{4}(b,x_{0})b_{\frac{1}{2}}y_{0}+((b_{\frac{1}{2}}x_{0})y_{0})b_{0}-(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})=0$
@@
$+\frac{1}{2}(b,x_{0})b_{\frac{1}{2}}y_{0}+2((b_{\frac{1}{2}}x_{0})y_{0})b_{0}-2(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})=0.$
@@
$-\Big(b_{\frac{1}{2}}x_{0})y_{0}-(b_{\frac{1}{2}}y_{0})x_{0}-b_{\frac{1}{2}}(x_{0}y_{0}\Big)-(b,x_{0}y_{0})b_{\frac{1}{2}},$
@@ 6. The proof of the identity
(1) By Lemma 4.4(2) and Theorem 5.1 we get,
@@
(2) Taking $x_{0}=\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}$ in (1), we get
@@
$=\Big[(1-(a,b))y_{\frac{1}{2}}-2y_{\frac{1}{2}}b_{0}\Big]\Big[\frac{1}{2}(1-(a,b))\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})b_{0}\Big]$
@@
$\textstyle{=\frac{1}{2}(1-(a,b))^{2}(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}-\frac{1}{8}(1-(a,b))^{2}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}}$
$\textstyle{+\frac{1}{4}(1-(a,b))(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}+\frac{1}{4}(1-(a,b))(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
@@
(3) Interchanging $x_{\frac{1}{2}},y_{\frac{1}{2}}$ in (2) adding and using Equation (4.1), we get
@@ 7. The proof of the identity
$2\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}-2\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}-\frac{1}{2}(b,x_{\frac{1}{2}}y_{0})b_{0}.$
@@
$+2\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}-2\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}-\frac{1}{2}(b,x_{\frac{1}{2}}y_{0})b_{0}.\qed$
@@
(7.6) $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}$
@@
$+(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}+(b,x_{\frac{1}{2}})y_{0}b_{0}+(b,x_{\frac{1}{2}}y_{0})b_{0}$
$=0.$
@@
$\textstyle{\frac{1}{2}z+2(b,y_{0})\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}+(b,y_{0})(b_{\frac{1}{2}}x_{\frac{1}{2}})b_{0}}$
@@
$\frac{1}{2}(1-(a,b))\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}\Big]$
@@
(3) By Lemma 6.2(1),
$(1-(a,b))\Big(b_{\frac{1}{2}}((y_{\frac{1}{2}}x_{0})b_{0})\Big)_{0}-2\Big(b_{\frac{1}{2}}(((y_{\frac{1}{2}}b_{0})x_{0})b_{0})$
@@
$=(1-(a,b))\Big(b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{0}b_{0}))\Big)_{0}$
$\textstyle{-\frac{1}{8}(b,x_{0})(b,y_{\frac{1}{2}})b_{0}}$
@@
By Proposition 7.4(2),
$4(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})-(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}$
$4\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}+2((a,b)-1)\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}\Big].$
@@
$((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0}))b_{0}=\frac{1}{2}(1-(a,b))\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}+\frac{1}{16}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}.$
@@
    Let $z:=\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-\Big((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}.$ If $(a,b)\neq 0,$ then $zb_{0}=\frac{1}{2}(1-(a,b)z.$
@@
$\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}+(a,b)\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+2((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})$
@@
But by Lemma 7.5(3), $2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))-2b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})-\frac{1}{2}(b,x_{\frac{1}{2}}y_{0})b_{0}=0.$ Hence, $z=2zb_{0}=(1-(a,b))z.$ It follows that $(a,b)z=0,$ so $z=0.$
@@ 8. The case $(a,b)=1$
Proof. (1) Suppose $b_{0}=0.$ Then, by Lemma 4.3(3), $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}=(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}-\Big(x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0})\Big)_{0}.$ By Lemma 4.13(2), $(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}=\frac{1}{2}(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}},$ because $(b,x_{\frac{1}{2}}y_{\frac{1}{2}})=(a,x_{\frac{1}{2}}y_{\frac{1}{2}}).$ So (1) follows from Theorem 3.7.
@@
$(3)\ \textstyle{b_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})=\frac{1}{2}x_{0}b_{0}.}$
@@
$\frac{1}{2}\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]+0+0$
$+0-\frac{1}{4}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.$
@@
$+(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}-\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.$
@@
$b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}=\\ \frac{1}{2}(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-3(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})+4\Big[b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))+b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))\Big],$ so (ii) follows from Lemma 8.2(6), since ${\rm char}(\mathbb{F})\neq 3.$
@@
    $+2\Big((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}+\frac{1}{2}(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}-2((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}=0.$
@@
$\textstyle{\frac{1}{4}(b,x_{\frac{1}{2}}y_{0})b_{0}+((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}-((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0})b_{0}+\frac{1}{8}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}}$
$\textstyle{+\frac{1}{8}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}-(b_{0},(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}=0}$
$\iff$
$\textstyle{\frac{1}{4}(b,x_{\frac{1}{2}}y_{0})b_{0}+((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}-((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0})b_{0}+\frac{1}{8}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}}$
$\textstyle{+\frac{1}{8}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}-\frac{1}{4}(b,y_{0})(b,x_{\frac{1}{2}})b_{0}=0.}$
@@
      $2(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})-\frac{1}{4}(b,x_{\frac{1}{2}})y_{0}b_{0}+(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})-((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0})b_{0}=0.$
@@
$\textstyle{4\Big[-\frac{1}{2}x_{\frac{1}{2}}b_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}\Big](b_{\frac{1}{2}}y_{0})}$
@@
$\textstyle{-2(x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})+\frac{1}{2}(b,x_{\frac{1}{2}})y_{0}b_{0}}$
@@
(ii) This follows from Lemma 8.4 and (i).

### v6

@@ 5. The proof that $b_{\frac{1}{2}}(x_{0}y_{0})=(b_{\frac{1}{2}}x_{0})y_{0}+(b_{\frac{1}{2}}y_{0})x_{0}$
$\textstyle{-\frac{1}{2}(b,y_{0})(a,b)b_{\frac{1}{2}}x_{0}-(b,y_{0})(b_{\frac{1}{2}}x_{0})b_{0}-(b,b_{\frac{1}{2}}x_{0})b_{\frac{1}{2}}y_{0}-(b,(b_{\frac{1}{2}}x_{0})y_{0})b_{\frac{1}{2}}+((b_{\frac{1}{2}}x_{0})y_{0})b_{0}}$
@@
$\textstyle{-\frac{1}{4}(b,x_{0})b_{\frac{1}{2}}y_{0}-\frac{1}{2}(b,y_{0})(a,b)b_{\frac{1}{2}}x_{0}-\frac{1}{4}(b,x_{0})(b,y_{0})b_{\frac{1}{2}}-(a,b)(b,x_{0})b_{\frac{1}{2}}y_{0}-(b,(b_{\frac{1}{2}}x_{0})y_{0})b_{\frac{1}{2}}}$
@@
$\textstyle{(2(a,b)-1)(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})+(b,x_{0})\Big(\frac{1}{2}(1-(a,b))b_{\frac{1}{2}}y_{0}+\frac{1}{4}(b,y_{0})b_{\frac{1}{2}}\Big)+2(a,b)(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})}$
$\textstyle{-\frac{1}{4}(b,x_{0})b_{\frac{1}{2}}y_{0}-\frac{1}{2}(b,y_{0})(a,b)b_{\frac{1}{2}}x_{0}-\frac{1}{4}(b,x_{0})(b,y_{0})b_{\frac{1}{2}}-(b,(b_{\frac{1}{2}}x_{0})y_{0})b_{\frac{1}{2}}}$
@@
$\textstyle{2(a,b)(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})+2(a,b)(b_{\frac{1}{2}}y_{0})(x_{0}b_{0})-\frac{1}{2}(a,b)(b,y_{0})b_{\frac{1}{2}}x_{0}-\frac{1}{2}(a,b)(b,x_{0})b_{\frac{1}{2}}y_{0}}$
$\textstyle{-\frac{1}{2}(a,b)(x_{0}b_{0},y_{0})b_{\frac{1}{2}}}$
$\textstyle{+\frac{1}{4}(b,x_{0})b_{\frac{1}{2}}y_{0}+((b_{\frac{1}{2}}x_{0})y_{0})b_{0}-(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})=0}$
@@
$\textstyle{+\frac{1}{2}(b,x_{0})b_{\frac{1}{2}}y_{0}+2((b_{\frac{1}{2}}x_{0})y_{0})b_{0}-2(b_{\frac{1}{2}}x_{0})(y_{0}b_{0})=0.}$
@@
$+\Big(b_{\frac{1}{2}}x_{0})y_{0}-(b_{\frac{1}{2}}y_{0})x_{0}-b_{\frac{1}{2}}(x_{0}y_{0}\Big)-(b,x_{0}y_{0})b_{\frac{1}{2}},$
@@ 6. The proof of the identity
(1) By Lemma 4.4(2) and Theorem 5.1. using also Theorem 4.9(4a), we get,
@@
(2) Taking $x_{0}=\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}$ in (1), and using $\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}z=(b_{\frac{1}{2}}x_{\frac{1}{2}})z-\frac{1}{4}(b,x_{\frac{1}{2}})z,$ for $z\in\{y_{\frac{1}{2}},y_{\frac{1}{2}}b_{0}\},$ and also Lemma 4.13(5), we get
@@
$=\textstyle{\Big[(1-(a,b))y_{\frac{1}{2}}-2y_{\frac{1}{2}}b_{0}\Big]\Big[\frac{1}{2}(1-(a,b))\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})b_{0}\Big]}$
@@
$\textstyle{=\frac{1}{2}(1-(a,b))^{2}(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}-\frac{1}{8}(1-(a,b))^{2}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+\frac{1}{4}(1-(a,b))(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
$\textstyle{-(1-(a,b))(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{\frac{1}{2}}b_{0})+\frac{1}{4}(1-(a,b))(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}}$
@@
(3) Interchanging $x_{\frac{1}{2}},y_{\frac{1}{2}}$ in (2) adding and using Equation (4.1) for $(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}})b_{0}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}})b_{0},$ we get
@@
$\textstyle{+\frac{1}{4}\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]}$
@@ 7. The proof of the identity
$\textstyle{2\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}-2\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}-\frac{1}{2}(b,x_{\frac{1}{2}}y_{0})b_{0}.}$
@@
$\textstyle{+2\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}-2\Big(b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})\Big)_{0}-\frac{1}{2}(b,x_{\frac{1}{2}}y_{0})b_{0}.}\qed$
@@
$+(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}+(b,x_{\frac{1}{2}})y_{0}b_{0}+(b,x_{\frac{1}{2}}y_{0})b_{0}=0.\qed$
@@
$\textstyle{\frac{1}{2}(1-(a,b))z+2(b,y_{0})\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}+(b,y_{0})(b_{\frac{1}{2}}x_{\frac{1}{2}})b_{0}}$
@@
(3) By Lemma 6.2(1) and part (2), using Theorem 4.9,
$(1-(a,b))\Big(b_{\frac{1}{2}}((y_{\frac{1}{2}}x_{0})b_{0})\Big)_{0}-2\Big(b_{\frac{1}{2}}(((y_{\frac{1}{2}}b_{0})x_{0})b_{0})\Big)_{0}$
@@
$=(1-(a,b))\Big(b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{0}b_{0}))\Big)_{0}+(1-(a,b))\Big((b_{\frac{1}{2}}x_{0})(y_{\frac{1}{2}}b_{0})\Big)_{0}$
$\textstyle{-\frac{1}{4}(1-(a,b))(b,y_{\frac{1}{2}})x_{0}b_{0}-\frac{1}{8}(b,x_{0})(b,y_{\frac{1}{2}})b_{0}}$
@@
By Proposition 7.4(2), $4(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})-(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}\quad=\\ 4\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))\Big)_{0}+2((a,b)-1)\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}\Big].$
@@
$m=\frac{1}{2}(1-(a,b))\Big((b_{\frac{1}{2}}y_{0})(x_{\frac{1}{2}}b_{0})\Big)_{0}+\frac{1}{16}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}.$
@@
    Let $z:=\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}.$ If $(a,b)\neq 0,$ then $zb_{0}=\frac{1}{2}(1-(a,b)z.$
@@
$\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}+(a,b)\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+2\Big((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}$
@@
But by Lemma 7.5(3), $2b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{0}b_{0}))-2b_{\frac{1}{2}}((x_{\frac{1}{2}}b_{0})y_{0})-\frac{1}{2}(b,x_{\frac{1}{2}}y_{0})b_{0}=0.$ Hence, by Proposition 7.8(2), $z=2zb_{0}=(1-(a,b))z.$ It follows that $(a,b)z=0,$ so $z=0.$
@@ 8. The case $(a,b)=1$
Proof. (1) Suppose $b_{0}=0.$ Then, by Lemma 4.3(3), $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}=(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}-\Big(x_{\frac{1}{2}}(b_{\frac{1}{2}}y_{0})\Big)_{0}.$ By Lemma 4.13(2), $(x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}=\frac{1}{2}(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}},$ because $(b,x_{\frac{1}{2}}y_{\frac{1}{2}})=(a,x_{\frac{1}{2}}y_{\frac{1}{2}}).$ So (1) follows from Theorem 5.1, and Theorem 3.7.
@@
$(3)\ \textstyle{b_{\frac{1}{2}}(b_{\frac{1}{2}}x_{0})=\frac{1}{2}(b,x_{0})a+\frac{1}{2}x_{0}b_{0}.}$
@@
$\textstyle{\frac{1}{2}\Big[(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}\Big]+0+0}$
$\textstyle{+0-\frac{1}{4}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
@@
$\textstyle{+(b,y_{\frac{1}{2}})x_{\frac{1}{2}}b_{0}+(b,x_{\frac{1}{2}})y_{\frac{1}{2}}b_{0}-\frac{1}{2}(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}.}$
@@
$b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{\frac{1}{2}})+(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{\frac{1}{2}}+(b_{\frac{1}{2}}y_{\frac{1}{2}})x_{\frac{1}{2}}=\\ \frac{1}{2}(b,y_{\frac{1}{2}})x_{\frac{1}{2}}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{\frac{1}{2}}+(a,x_{\frac{1}{2}}y_{\frac{1}{2}})b_{\frac{1}{2}}-3(b_{0},x_{\frac{1}{2}}y_{\frac{1}{2}})+4\Big[b_{\frac{1}{2}}(y_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0}))+b_{\frac{1}{2}}(x_{\frac{1}{2}}(y_{\frac{1}{2}}b_{0}))\Big],$ so (2i) follows from Lemma 8.2(6), since ${\rm char}(\mathbb{F})\neq 3.$
@@
    $\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}y_{0})\Big)_{0}+\Big((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}\Big)_{0}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}+\\ 2\Big((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}+\frac{1}{2}(b,y_{0})\Big(b_{\frac{1}{2}}x_{\frac{1}{2}}\Big)_{0}-2((b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}=0.$
@@
$\Big[\Big(b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})\Big)_{0}+(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}}-(b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0}\Big]b_{0}$
$\textstyle{+\frac{1}{8}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}+\frac{1}{8}(b,x_{\frac{1}{2}})(b,y_{0})b_{0}-(b_{0},(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})b_{0}=0.}$
Since $(b_{0},(b_{\frac{1}{2}}y_{0})x_{\frac{1}{2}})=\frac{1}{4}(b,x_{\frac{1}{2}})(b,y_{0})b_{0},$ and $b_{\frac{1}{2}}(x_{\frac{1}{2}}b_{0})b_{0}=\frac{1}{4}(b,x_{\frac{1}{2}}y_{0})b_{0},$ (7i) holds.
@@
      $2\Big((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}-\frac{1}{4}(b,x_{\frac{1}{2}})y_{0}b_{0}+(b_{\frac{1}{2}}x_{\frac{1}{2}})(y_{0}b_{0})-((b_{\frac{1}{2}}x_{\frac{1}{2}})y_{0})b_{0}=0.$
@@
$\textstyle{4\Big(\Big[-\frac{1}{2}x_{\frac{1}{2}}b_{0}+\frac{1}{4}(b,x_{\frac{1}{2}})b_{\frac{1}{2}}\Big](b_{\frac{1}{2}}y_{0})\Big)_{0}}$
@@
$\textstyle{-2\Big((x_{\frac{1}{2}}b_{0})(b_{\frac{1}{2}}y_{0})\Big)_{0}+\frac{1}{2}(b,x_{\frac{1}{2}})y_{0}b_{0}}$
@@
(ii) Computing 22 times the equality in (i) minus the equality in Lemma 8.4 yields (ii).